(8) (14 pts). Solve y + y = cos(t) +e-t, y(0) = y(0) = 0.
Answer #1

Som y + y = cast + ét , y0) = y(0) = 0 • Ag + 4 = as t e ie (²+1) y = cost + et, where D = d The auxiliary equation is 0+1=0the general sol is 94) = yelt) + YpCt) je ya) = G ast + 6 sint + sint + + ét -0 y(t) = 4, sint + G cast + 1 [t cast + eint)